OpenStudy (anonymous):

Tell whether the linear system has zero, one or more solution 4x-3y=6 8x=6y+10

OpenStudy (cwrw238):

first rewrite them to make comparing them easier 4x - 3y = 6 8x - 6y = 10

OpenStudy (cwrw238):

now multiply first equation by 2:- 8x - 6y = 12 and second equation is 8x - 6y = 10 do you notice anything?

OpenStudy (anonymous):

Yeah they're the same just have different answers

OpenStudy (cwrw238):

yes left sides are the same there are no solutions since left side cannot be equal to both 10 and 12

OpenStudy (anonymous):

@Thank you!

OpenStudy (cwrw238):

yw